OpenStudy (anonymous):

Help please!! I'm failing Math!! I really need help! A test consists of multiple choice questions worth 3 points and essay questions worth 5 points. The test has 17 questions that are worth a total of 65 points. Use a model to determine how many multiple choice questions are on the test.

OpenStudy (anonymous):

i dont get how you are supposed to figure it out

OpenStudy (anonymous):

in a normal test the maximum essays you get is 5 so i would say there's 12 multiple choice questions im not sure ( not using math just my head) lol

OpenStudy (anonymous):

can anyone explain this please?

OpenStudy (anonymous):

ok listen

OpenStudy (anonymous):

you make 2 equations here let multiple choice be x and essay be y then there are 17 questions right? so x+y =17 and x is for 3 marks and y is for 5 total marks are 65 so 3x+5y =65 so now you have 2 equations x+y=17 and 3x+5y=65 solve them you will get the answer

OpenStudy (anonymous):

ok thank you!!

OpenStudy (anonymous):

did you understand?

OpenStudy (anonymous):

yeah :)